OpenStudy (anonymous):

8 = 33 + y what is the value of y? please help

OpenStudy (anonymous):

-25

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

8 = 33 + y 8 - 33 = 33 - 33 + y